# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys using DFPT

### v3

@@ Acknowledgements
This work was supported in part by NASA EPSCoR Grant # 80NSSC21M0171. We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing computational resources and system support for this work.
@@
The following figures provide additional computational results for binaries not discussed in the main text. Figures 912 illustrate the spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for MoTa, MoW, NbTa and WNb, as a function of composition. Figures 1316 illustrate the relaxation field of $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$ as a function of composition for MoNb, MoTa, MoW and WNb.
@@
Figure 9: Spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for: (a)-(c)  $\rm{Mo_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}Ta_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 10: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Mo_{0.25}W_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}W_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}W_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 11: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Nb_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Nb_{0.5}Ta_{0.5}}$;  and (g)-(h) $\rm{Nb_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 12: Spatial dependence of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{W_{0.25}Nb_{0.75}}$;  (d)-(f) $\rm{W_{0.5}Nb_{0.5}}$; and (g)-(h) $\rm{W_{0.75}Nb_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 13: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.5}Nb_{0.5}}$, and $\rm{Mo_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 14: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, and $\rm{Mo_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 15: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}W_{0.75}}$, $\rm{Mo_{0.5}W_{0.5}}$, and $\rm{Mo_{0.75}W_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 16: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Nb_{0.75}}$, $\rm{W_{0.5}Nb_{0.5}}$, and $\rm{W_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.

### v4

@@ Acknowledgements
This work was supported in part by NASA EPSCoR Grant # 80NSSC21M0171. We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing computational resources and system support for this work. Comments and suggestions by anonymous reviewers are gratefully acknowledged.
@@
A.1 Derivation of nuclear relaxation expressions for Young’s modulus $E$ and Poisson ratio $\nu$.
Young’s modulus $E$. Starting from Eq. (25), and the Voigt expression for $E$ (Eq. (12)), we want to derive an expression for $E_{\text{nuclear}}$ in terms of the quantities $\{B_{n},G_{n},B_{r},G_{r}\}$, where $B_{n}=B_{\text{nuclear}}$, $G_{n}=G_{\text{nuclear}}$, $B_{r}=B_{\text{rigid}}$, and $G_{r}=G_{\text{rigid}}$ as defined in the main text.
We have:
$E_{\text{relaxed}}=E_{\text{Voigt}}$ $=\frac{9B_{\text{Voigt}}G_{\text{Voigt}}}{3B_{\text{Voigt}}+G_{\text{Voigt}}}=\frac{9\left(B_{r}-B_{n}\right)\left(G_{r}-G_{n}\right)}{3\left(B_{r}-B_{n}\right)+\left(G_{r}-G_{n}\right)}$
$=\frac{9\left(B_{r}G_{r}-B_{r}G_{n}-B_{n}G_{r}+B_{n}G_{n}\right)}{\left(3B_{r}+G_{r}\right)-\left(3B_{n}+G_{n}\right)}=\frac{9B_{r}G_{r}-9\left(B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}\right)}{\left(3B_{r}+G_{r}\right)\left(1-E_{1}\right)},$
where we have defined
$E_{1}\equiv\frac{3B_{n}+G_{n}}{3B_{r}+G_{r}}.$ (A.1)
If the nuclear relaxation contributions to $B$ and $G$ are assumed to be small compared to the rigid values for $B$ and $G$, then $E_{1}\ll 1$; this condition will be shown to be an excellent approximation in Section A.2 below. We can therefore approximate $1/(1-E_{1})=(1-E_{1})^{-1}$ by its first order Taylor series value, $1+E_{1}$. Noting that
$E_{\text{rigid}}=\frac{9B_{r}G_{r}}{3B_{r}+G_{r}},$
and defining
$E_{2}\equiv\frac{B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}}{B_{r}G_{r}},$ (A.2)
we have:
$E_{\text{Voigt}}$ $=\frac{9B_{r}G_{r}-9\left(B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}\right)}{\left(3B_{r}+G_{r}\right)\left(1-E_{1}\right)}$
$\approx\left(\frac{9B_{r}G_{r}}{3B_{r}+G_{r}}-\frac{9\left(B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}\right)}{3B_{r}+G_{r}}\right)\left(1+E_{1}\right)$
$=\frac{9B_{r}G_{r}}{3B_{r}+G_{r}}\left(1-\frac{\left(B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}\right)}{B_{r}G_{r}}\right)\left(1+E_{1}\right)$
$=E_{\text{rigid}}(1-E_{2})(1+E_{1})=E_{\text{rigid}}(1-E_{2}+E_{1}-E_{2}E_{1})$
$=E_{\text{rigid}}-E_{\text{rigid}}(E_{2}-E_{1}+E_{2}E_{1}).$
Thus,
$E_{\text{nuclear}}\approx E_{\text{rigid}}(E_{2}-E_{1}+E_{2}E_{1}).$ (A.3)
Possion’s ratio $\nu$. Starting from Eq. (30):
$\nu_{\text{relaxed}}=\nu_{\text{Voigt}}$ $=\frac{3B_{\text{Voigt}}-2G_{\text{Voigt}}}{2\left(3B_{\text{Voigt}}+G_{\text{Voigt}}\right)}=\frac{3\left(B_{r}-B_{n}\right)-2\left(G_{r}-G_{n}\right)}{2\left(3\left(B_{r}-B_{n}\right)+\left(G_{r}-G_{n}\right)\right)}$
$=\frac{\left(3B_{r}-2G_{r}\right)-\left(3B_{n}-2G_{n}\right)}{2\left(3B_{r}+G_{r}\right)-2\left(3B_{n}+G_{n}\right)}=\frac{\left(3B_{r}-2G_{r}\right)-\left(3B_{n}-2G_{n}\right)}{2\left(3B_{r}+G_{r}\right)\left(1-\nu_{1}\right)},$
where we have defined
$\nu_{1}=\frac{3B_{n}+G_{n}}{3B_{r}+G_{r}}.$ (A.4)
Note that $\nu_{1}=E_{1}$ from the Young’s modulus derivation. We again assume that $\nu_{1}\ll 1$, so that $1/(1-\nu_{1})\approx 1+\nu_{1}$. This gives:
$\nu_{\text{Voigt}}$ $\approx\left(\frac{\left(3B_{r}-2G_{r}\right)-\left(3B_{n}-2G_{n}\right)}{2\left(3B_{r}+G_{r}\right)}\right)(1+\nu_{1})=\left(\frac{3B_{r}-2G_{r}}{2(3B_{r}+G_{r})}-\frac{(3B_{n}-2G_{n})}{2(3B_{r}+G_{r})}\right)(1+\nu_{1})$
$=\nu_{\text{rigid}}\left(1-\nu_{2}\right)(1+\nu_{1})=\nu_{\text{rigid}}\left(1-\nu_{2}+\nu_{1}-\nu_{2}\nu_{1}\right)$
$=\nu_{\text{rigid}}-\nu_{\text{rigid}}\left(\nu_{2}-\nu_{1}+\nu_{2}\nu_{1}\right),$
where we have defined
$\nu_{2}=\frac{3B_{n}-2G_{n}}{3B_{r}-2G_{r}}.$ (A.5)
Thus,
$\nu_{\text{nuclear}}\approx\nu_{\text{rigid}}\left(\nu_{2}-\nu_{1}+\nu_{2}\nu_{1}\right).$ (A.6)
A.2 Nuclear relaxation field validation tables
Table 3: Comparison of Abinit-computed nuclear-relaxation values ($E_{nuclear}$) with summation of approximated atomic nuclear relaxation values ($\sum_{p}E_{p,nuclear}$) over all atoms $p$ in the supercell. For $E_{1}$ the max value corresponds to maximum value of $E_{1}$ defined in Eq.( A.1) across all atoms in the supercell.
Composition
$E_{1}$
(max value)
$E_{nuclear}$ $\sum_{p}E_{p,nuclear}$ Composition
$E_{1}$
(max value)
$E_{nuclear}$ $\sum_{p}E_{p,nuclear}$
AB at.% of A AB at.% of A
MoNb 25 8.16E-04 2.7267 2.7253 MoTa 25 1.07E-03 9.2285 9.1913
50 5.24E-04 1.9929 1.9933 50 5.19E-04 5.1921 5.1829
75 3.30E-04 1.0192 1.0193 75 2.48E-04 1.3723 1.3718
MoW 25 6.91E-05 0.2084 0.2084 NbTa 25 1.22E-04 1.7917 1.7898
50 1.00E-04 0.3951 0.3951 50 2.57E-04 4.3686 4.3573
75 6.98E-05 0.1786 0.1786 75 1.86E-04 2.6172 2.6130
WNb 25 9.75E-04 1.1373 1.1384 WTa 25 9.98E-04 4.3394 4.3367
50 9.57E-04 2.1130 2.1150 50 7.37E-04 4.1908 4.1907
75 4.21E-04 1.0581 1.0584 75 3.24E-04 1.0794 1.0796
Table 4: Comparison of Abinit-computed nuclear-relaxation values ($\nu_{nuclear}$) with summation of approximated atomic nuclear relaxation values ($\sum_{p}\nu_{p,nuclear}$) over all atoms $p$ in the supercell. For $\nu_{1}$ the max value corresponds to maximum value of $\nu_{1}$ defined in Eq.( A.4) across all atoms in the supercell.
Composition
$\nu_{1}$
(max value)
$\nu_{nuclear}$ $\sum_{p}\nu_{p,nuclear}$ Composition
$\nu_{1}$
(max value)
$\nu_{nuclear}$ $\sum_{p}\nu_{p,nuclear}$
AB at.% of A AB at.% of A
MoNb 25 8.16E-04 -1.616E-03 -1.603E-03 MoTa 25 1.07E-03 -6.716E-03 -6.652E-03
50 5.24E-04 -5.552E-04 -5.510E-04 50 5.19E-04 -3.175E-03 -3.157E-03
75 3.30E-04 -2.688E-04 -2.680E-04 75 2.48E-04 -7.135E-04 -7.123E-04
MoW 25 6.91E-05 3.741E-05 3.739E-05 NbTa 25 1.22E-04 -1.560E-03 -1.558E-03
50 1.00E-04 7.868E-05 7.861E-05 50 2.57E-04 -3.952E-03 -3.939E-03
75 6.98E-05 4.592E-05 4.590E-05 75 1.86E-04 -2.465E-03 -2.459E-03
WNb 25 9.57E-04 4.534E-04 4.520E-04 WTa 25 9.75E-04 4.534E-04 4.520E-04
50 9.57E-04 8.572E-04 8.501E-04 50 7.37E-04 -1.290E-03 -1.276E-03
75 4.21E-04 2.917E-04 2.909E-04 75 3.24E-04 -6.232E-05 -6.192E-05
A.3 Binary alloy anisotropy plots and nuclear relaxation fields
@@
The following figures provide additional computational results for binaries not discussed in the main text. Figures 1114 illustrate the spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for MoTa, MoW, NbTa and WNb, as a function of composition. Figures 1522 illustrate the nuclear-relaxation fields for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$, $E$ and $\nu$ as a function of composition for MoNb, MoTa, MoW and WNb.
@@
Figure 11: Anisotropy plots of $G$ (GPa), $E$ (GPa), and $\nu$ for: (a)-(c)  $\rm{Mo_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}Ta_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 12: Anisotropy plots of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Mo_{0.25}W_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}W_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}W_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 13: Anisotropy plots of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{Nb_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{Nb_{0.5}Ta_{0.5}}$;  and (g)-(h) $\rm{Nb_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 14: Anisotropy plots of $G$ (GPa), $E$ (GPa) and $\nu$ for: (a)-(c) $\rm{W_{0.25}Nb_{0.75}}$;  (d)-(f) $\rm{W_{0.5}Nb_{0.5}}$; and (g)-(h) $\rm{W_{0.75}Nb_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 15: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.5}Nb_{0.5}}$, and $\rm{Mo_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 16: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.5}Nb_{0.5}}$, and $\rm{Mo_{0.75}Nb_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 17: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, and $\rm{Mo_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 18: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, and $\rm{Mo_{0.75}Ta_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 19: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}W_{0.75}}$, $\rm{Mo_{0.5}W_{0.5}}$, and $\rm{Mo_{0.75}W_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 20: Relaxation field (values in GPa) constructed as described in text for $\rm{Mo_{0.25}W_{0.75}}$, $\rm{Mo_{0.5}W_{0.5}}$, and $\rm{Mo_{0.75}W_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 21: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Nb_{0.75}}$, $\rm{W_{0.5}Nb_{0.5}}$, and $\rm{W_{0.75}Nb_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 22: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Nb_{0.75}}$, $\rm{W_{0.5}Nb_{0.5}}$, and $\rm{W_{0.75}Nb_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.